OpenStudy (anonymous):

Derivative of 2^(-1/x)

Parth (parthkohli):

\( \color{Black}{\Rightarrow x^n \Rightarrow nx^{n - 1} }\)

Parth (parthkohli):

This is how you find the derivative ^

OpenStudy (lgbasallote):

that only works when x is base and exponent is constant

OpenStudy (lgbasallote):

derivative of \(a^x = a^x\ln a\) if i remember right

OpenStudy (anonymous):

how might I end up with the answer 2^(-1/x) ln2/x^2

OpenStudy (lgbasallote):

like i said derivative of a^x = a^x ln a here your a is 2 so it's 2^(-1/x) ln 2 but wait! you have to take the derivative of -1/x as well

OpenStudy (lgbasallote):

derivative of -1/x = 1/x^2

OpenStudy (lgbasallote):

so combining all of that you will get 2^(-1/x) ln 2 / x^2 got it?

OpenStudy (anonymous):